I need some help with this problem?
The​ half-life of radium is 1690 years. If 30 grams are present​ now, how much will be present in 210 ​years?

Answer by ikleyn(52781)

.

The general formula for the remaining amount is


    M(t) = M(0)*2^(-t/1690),


where  M(0)  is the initial mass in grams;  t is the time in years.


Substitute your data to get


    M(in 210 years) = 30*2^(-210/1690) = 30*2^(-0.12426) = 27.524 grams (approximately)    ANSWER

Solved.